Determine the pH (to two decimal places) of a solution prepared by adding 239 mL of 0.750 M sodium dihydrogen citrate (NaC3H5O(COOH)2COO) to 3.99 g of sodium hydrogen citrate (Na2C3H5O(COOH)(COO)2).

Assume the volume of the solution does not change and that the 5% approximation is valid. pKa=4.77
Answer #1

we have

pKa = 4.77

moles of sodium hydrogen citrate = 3.99 g / 236.09 g/mol = 0.0169 mol

moles of sodium dihydrogen citrate = 239 ml x 0.750 M = 0.17925 mol

therefore

using the henderson equation , we have

pH = pKa + log[sodium hydrogen citrate]/[sodium dihydrogen citrate]

= 4.77 + log( 0.0169 / 0.17925 )

= 4.77 + log(0.0943)

= 4.77 - 1.03

= 3.74

therefore pH of the solution is = 3.74